# Beyond Image Prior: Embedding Noise Prior into Conditional Denoising Transformer

### v1

@@ 4.3 Experiments on Image Denoising
TABLE II: Quantitative comparisons of different real image denoising methods on several benchmarks. Particularly, the criteria of SIDD and DND benchmarks are obtained from the corresponding online server.
[…]
Noisy - - 23.66 0.4848 29.56 0.3347 29.84 0.7015
RIDNet [3] ICCV’19 1.5 38.77 0.9511 38.98 0.9076 39.24 0.9513
VDN [16] NeurIPS’19 7.8 39.36 0.9562 39.49 0.9117 39.30 0.9493
MIRNet [81] ECCV’20 31.8 39.72 0.9586 39.80 0.9147 39.88 0.9543
MPRNet [5] CVPR’21 15.7 39.71 0.9586 39.80 0.9149 39.82 0.9540
Uformer [8] CVPR’22 50.8 39.89 0.9594 39.97 0.9160 40.05 0.9562
Restormer [9] CVPR’22 26.1 40.02 0.9603 40.09 0.9171 40.03 0.9564
NAFNet [82] ECCV’22 29.1 39.97 0.9599 40.04 0.9166 39.10 0.9495
MSANet [83] NeurIPS’22 8.6 39.56 0.9575 39.70 0.9131 39.65 0.9553
CAT [49] NeurIPS’22 25.8 40.01 0.9600 40.09 0.9167 40.05 0.9561
MIRNetv2 [4] TPAMI’23 5.9 39.84 0.9593 39.91 0.9154 39.86 0.9550
ShuffleFormer [50] ICML’23 50.1 40.00 0.9603 40.08 0.9168 40.01 0.9560
GRL [52] CVPR’23 19.8 39.89 0.9595 40.01 0.9161 39.76 0.9540
ART [51] ICLR’23 25.7 39.96 0.9598 40.03 0.9164 40.05 0.9557
VIRNet [28] TPAMI’24 15.4 39.70 0.9586 39.78 0.9148 39.77 0.9533
MambaIR [47] ECCV’24 23.2 39.89 0.9598 39.97 0.9164 39.83 0.9542
Condformer Ours 27.0 40.21 0.9612 40.23 0.9176 40.10 0.9562
@@ 4.3.1 On synthetic images
In this section, we conduct Poisson-Gaussian image denoising experiments on synthetic validation datasets. Note that, recent blind denoisers were commonly trained on noise models as Gaussian (G), spatially-variant Gaussian (SV-G) and Poisson-Gaussian (P-G) distributions. We formulate the corresponding noises as follows,
@@
obviously, Gaussian noise model is a specific case of P-G noise model when $\sigma_{\text{s}}=0$, and SV-G noise model is more fine-grain as applying a spatial map $\mathbf{M}$ to represent the variance of Gaussian, thus P-G noise model can be regarded as a specific case of SV-G by setting the map relative to image context $\mathbf{L}$.
In comparison of synthetic image denoising, several current state-of-the-art blind denoisers are selected, including: 1) Gaussian noise model drived: DnCNN [15] and Restormer [9]; 2) SV-G noise model drived: VDN [16], DRANet [77] and VIRNet [28]; 3) P-G noise model drived: FBI-D [78], VBDNet [79], CLIPDenoising [80] and our Condformer. Moreover, we re-train the Restormer [9] and MambaIR [47] denoisers on P-G noise model with same settings as our Condformer for fair comparisons.
As reported in TABLE I, quantitative comparisons on three public validation datasets with various noise levels show that, our Condformer achieves superior performances against other methods. On denoising accuracy, our Condformer obtains higher PSNR values especially when $\sigma_{\text{s}}>0$ and shows slight disadvantages compared with the Restormer trained on Gaussian noise model when $\sigma_{\text{s}}=0$. On model generalization, under same experimental settings, our Condformer achieves excellent performances on all noise levels, instead Restormer and MambaIR show lower generalization on various noise levels. That is, our Condformer possesses a complete and conditional optimization space for divide-and-conquer, which is naturally stronger than any unconditional models. Furthermore, the qualitative comparisons shown in Fig. 8 demonstrate our Condformer can handle any degree of noisy corruption, being capable of detail preservation and noise removal well.
@@ 4.3.2 On real images
In this section, we demonstrate the effectiveness of the proposed Condformer for real image denoising. We compare our Condformer with several state-of-the-art real image denoising methods on SIDD validation, SIDD benchmark, and DND benchmark datasets. We select several representative CNN-based denoisers and current state-of-the-art Transformer-based and Mamba-based denoisers, including: 1) CNN-based denoiser: RIDNet [3], VDN [16],MIRNet [81], MPRNet [5], NAFNet [82], MSANet [83], MIRNetv2 [4] and VIRNet [28]; 2) Transformer-based denoiser: Uformer [8], Restormer [9], CAT [49], ShuffleFormer [50], GRL [52] and ART [51]; 3) Mamba-based denoiser: MambaIR [47]. As reported in TABLE II, our Condformer achieves the highest PSNR and SSIM criteria over all the compared denoising methods. Note that, since the clean groundtruth images of SIDD and DND benchmarks are inaccessible, we upload the denoised results of all the considered methods on their online servers for testing.
@@
Qualitatively on the in-distribution SIDD validation and benchmark datasets, as shown in Fig. 9 and Fig. 10, our Condformer can not only successfully remove the noises but also keep the details well, including textures, edges and natural structures. After conditional optimization with noise prior, our Condformer could handle each noisy image in a sensor-specific way, where the denoising intensity is proportional to the corruption degree of the noisy observation. Especially under low-light imaging environments, noises prone to overwhelm the image contexts as introducing higher ISO level and indirectly enlarging the shot noises. Furthermore, Fig. 11 and Fig. 12 shows that our Condformer has excellent generalization ability as handling the out-of-distribution DND noisy images well, on both detail preservation and undesired noise removal. Note that, since the images have low visibility under low-light imaging environment, we employ image normalization on each denoised result for better visualization.
TABLE III: Quantitative comparisons of different Transformer-based denoising methods on computational complexity. The FLOPs, Memory and Time are calculated when processing a single $512\times 512\times 3$ sRGB image on a NVIDIA A800 GPU.
@@
Uformer [8] 50.8 343.1 2.89 0.21
Restormer [9] 26.1 564.0 3.81 0.36
CAT [49] 25.8 543.5 10.83 15.76
ShuffleFormer [50] 50.5 344.4 2.97 0.23
GRL [52] 19.8 5012.3 12.38 10.45
ART [51] 25.7 542.8 3.66 0.41
@@ 4.3.3 On computational complexity
In this section, we mainly conduct a comparison of our Condformer and other Transformer-based denoising methods on several computational complexity criteria. In detail, the number of parameters represents the model size for transmission and storage necessaries, FLOPs and Time indicate the time complexity of model, and Memory is the memory usage when running model on a GPU device. The last three criteria are calculated when processing a single $512\times 512\times 3$ sRGB image on a NVIDIA A800 GPU. To avoid randomness, the running time is averaged on handling 100 images with torch.cuda.Event timer. As reported in TABLE III, our Condformer shows relatively trade-off on the complexities, which is a resource-friendly method.
@@ 4.4 Model Analysis
TABLE IV: Ablation study of Condformer with different priors on SIDD Validation dataset.
@@
(M) (G) (s) (dB)
- - 26.11 563.96 0.361 39.68
[…]
$(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ 26.74 564.22 0.368 39.81
$(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ 27.02 565.35 0.368 39.90
$(\sigma_{\text{s}},\sigma_{\text{r}})$ 26.41 565.17 1.012 39.90
@@ 4.4.1 Investigation on Condformer
Aiming at guiding the model to learn from image prior and noise prior separately, the core of our Condformer is embedding the noise prior effectively in the latent space for conditional optimization. Thus, we conduct an ablation study of Condformer in TABLE IV, note that all the models in this section are trained during 100K iterations, where the image size is $3\times 128\times 128$ and batch size is set to 8 for training on a single NVIDIA A800 GPU.
Since the key of our Condformer is constructing CondSA blocks in the latent module in Restormer, we first train a baseline model which is same sa Restormer and achieves PSNR of 39.68dB on SIDD validation dataset. Subsequently, to avoid deviation on the performance as introducing an auxiliary vector, we firstly build a void CondSA module by embedding zero vector, and find it gain nothing. However, after embedding the estimated noise prior $(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ from LoNPE network by a naive concatenation with the input tensor of CondSA, the corresponding model achieves about 0.1 PSNR gain, demonstrating the positive role of noise prior. Furthermore, to learn an effective representation of the correlation between-in the intermediate image features and noise prior, a LFM module is designed for correlation representation of noise prior and image features. The result shows that an effective fusion module can significantly boost the contribution of the noise prior. As reported in TABLE IV, the fourth model achieves higher performances than the afore three models, especially gains more than 0.2dB of PSNR against the baseline.
@@
TABLE V: Investigation on LoNPE algorithm with different patch-sampling hyperparameters. Particularly, $\mathcal{O}$, $m/n$ and $\lambda_{\mathcal{S}}$ denote the sampling size, ratio and index, respectively.
[…]
$\mathcal{O}$ $m/n$ $\lambda_{\mathcal{S}}$ $\Delta(\sigma_{\text{s}})$ $\Delta(\sigma_{\text{r}})$
$8\times 8$ $5$% 0.094 0.056 0.31
$8\times 8$ $5$% 0.052 0.037 0.36
$8\times 8$ $10$% 0.052 0.037 0.36
[…]
$16\times 16$ $20$% 0.029 0.022 1.13
[…]
LoNPE Network 0.032 0.026 $<0.01$
@@ 4.4.2 Investigation on LoNPE
We conduct an investigation on the effects of sampling size $\mathcal{O}$ and ratio $m/n$ in TABLE V. Since there is no groundtruth for noise prior in SIDD-Medium raw-domain training dataset, we synthesize noisy samples by randomly sampling the noises from a Poisson-Gaussian distribution with $(\sigma_{\text{s}},\sigma_{\text{r}})=(0.2,0.01)$ and adding it on a clean raw noisy image. The average error (i.e., mean absolute percentage error $\Delta(\cdot)$ [84]) and running time are reported in TABLE V, with the increase of patch size $\mathcal{O}$ and sampling ratio $m/n$, it might introduce a negative effect of image prior. In a word, a larger patch has more precise statistical values but might contains more image priors interfering the noise prior estimation, such as edges and textures. Empirically, we choose $\mathcal{O}=16\times 16$ and $m/n=10$% which provides more precise estimation of noise prior.
Besides, to further eliminate the image prior in each patch, we introduce a local smoothness criterion $\lambda_{\mathcal{S}}$ to filter the smooth patches from the original patch pools $\{I_{i}\}^{n}_{i=1}$, which can effectively boost the accuracy of noise prior estimation as reported in TABLE V. To further demonstrate the effectiveness of local smoothness criterion, we record the sampled local patches in Fig. 13. It is obvious that random sampling local patches from the raw image inevitably captures the numerous high-frequency details, which would interfere the estimation of noise prior because of the nonnegligible deviation on calculating the local luminance in Eq.(9). Instead, by employing the local smoothness criterion $\lambda_{\mathcal{S}}$, it is easy to filter the smooth patches from the original patch pools, which plays significant role on eliminating the image prior and help estimating noise prior.

### v2

@@ 4.3 Experiments on Image Denoising
In this section, we mainly conduct quantitatively and qualitatively experimental study on synthetic and real blind image denoising performance of the proposed Condformer.
@@ 4.3.1 On synthetic images
Due to the agnosticism of real noise prior, it is hard to conduct a comprehensive validation on image denoising under various noise priors. We firstly conduct Poisson-Gaussian blind image denoising experiments on synthetic validation datasets. We mainly synthesize three levels of shot noise prior $\sigma_{\text{s}}\in[0,0.15,0.3]$, and sample read noise priors $\sigma_{\text{r}}\in[15,25,50]/255$ with each fixed $\sigma_{\text{s}}$. Particularly, this setting follows the convention of mainstream Gaussian blind image denoising researches as $\sigma_{\text{s}}=0$ and $\sigma_{\text{r}}\in[15,25,50]/255$.
Recent blind denoisers were commonly trained on noise models as Gaussian (G), spatially-variant Gaussian (SV-G) and Poisson-Gaussian (P-G) distributions. We formulate the corresponding noises as follows,
@@
obviously, Gaussian noise model is a specific case of P-G noise model when $\sigma_{\text{s}}=0$, and SV-G noise model is more fine-grain as applying a spatial map $\mathbf{M}$ to represent the variance of Gaussian in pixel-wise level, thus P-G noise model can be regarded as a specific case of SV-G by setting the map relative to image context $\mathbf{L}$.
Table 3: Quantitative comparisons of different real image denoising methods on several benchmarks. Particularly, the criteria of SIDD and DND benchmarks are obtained from the corresponding online server.
[…]
Noisy - 23.66 0.4848 29.56 0.3347 29.84 0.7015
RIDNet 4 1.5 38.77 0.9511 38.98 0.9076 39.24 0.9513
VDN 63 7.8 39.36 0.9562 39.49 0.9117 39.30 0.9493
MIRNet 65 31.8 39.72 0.9586 39.80 0.9147 39.88 0.9543
MPRNet 66 15.7 39.71 0.9586 39.80 0.9149 39.82 0.9540
Uformer 56 50.8 39.89 0.9594 39.97 0.9160 40.05 0.9562
Restormer 67 26.1 40.02 0.9603 40.09 0.9171 40.03 0.9564
NAFNet 12 29.1 39.97 0.9599 40.04 0.9166 39.10 0.9495
MSANet 20 8.6 39.56 0.9575 39.70 0.9131 39.65 0.9553
CAT 76 25.8 40.01 0.9600 40.09 0.9167 40.05 0.9561
MIRNetv2 68 5.9 39.84 0.9593 39.91 0.9154 39.86 0.9550
ShuffleFormer 60 50.1 40.00 0.9603 40.08 0.9168 40.01 0.9560
GRL 30 19.8 39.89 0.9595 40.01 0.9161 39.76 0.9540
ART 70 25.7 39.96 0.9598 40.03 0.9164 40.05 0.9557
VIRNet 64 15.4 39.70 0.9586 39.78 0.9148 39.77 0.9533
MambaIR 21 23.2 39.89 0.9598 39.97 0.9164 39.83 0.9542
Condformer (Ours) 27.0 40.21 0.9612 40.23 0.9176 40.10 0.9562
In comparison of synthetic image denoising, several current state-of-the-art blind denoisers are selected, including:
1) Gaussian noise model drived: DnCNN 71 and Restormer 67;
2) SV-G noise model drived: VDN 63, DRANet 59 and VIRNet 64;
3) P-G noise model drived: FBI-Denoiser 7, VBDNet 31, CLIPDenoising 13 and our Condformer.
In addition, we re-train the Restormer 67 and MambaIR 21 denoisers on P-G noise model with same settings as our Condformer for fair comparisons, marked as Restormer and MambaIR.
As reported in Table 2, quantitative comparisons on three public validation datasets with various noise levels show that, our Condformer achieves superior performances against other methods. On denoising accuracy, our Condformer obtains higher PSNR values especially when handling Poisson-Gaussian noises ($\sigma_{\text{s}}>0$) and shows slight disadvantages compared with the state-of-the-art Restormer trained on Gaussian noise model when handing Gaussian noises ($\sigma_{\text{s}}=0$). On model generalization, under same experimental settings, our Condformer achieves excellent performances on all noise levels, instead Restormer and MambaIR show lower generalization on various noise levels. That is, our Condformer possesses a complete and conditional optimization space for divide-and-conquer, which is naturally stronger than any unconditional models.
Furthermore, we also provide the qualitative visual comparisons of restoring the Poisson-Gaussian noisy images in Fig. 8. Our Condformer can handle any degree of noisy corruptions, being capable of detail preservation and noise removal well. Especially, under severe corruptions, only our Condformer simultaneously recovers the details (such as textures of the helmet, edges of the windows), alleviates the distortions in smooth region and improves the fidelity of color.
@@ 4.3.2 On real images
In this section, we demonstrate the effectiveness of the proposed Condformer for real image denoising. We compare our Condformer with several state-of-the-art real image denoising methods on SIDD validation, SIDD benchmark, and DND benchmark datasets. We select several representative CNN-based denoisers and current state-of-the-art Transformer-based and Mamba-based denoisers, including:
1) CNN-based denoiser: RIDNet 4, VDN 63,MIRNet 65, MPRNet 66, NAFNet 12, MSANet 20, MIRNetv2 68 and VIRNet 64;
2) Transformer-based denoiser: Uformer 56, Restormer 67, CAT 76, ShuffleFormer 60, GRL 30 and ART 70;
3) Mamba-based denoiser: MambaIR 21.
As reported in Table 3, our Condformer achieves the highest PSNR and SSIM criteria over all the compared denoising methods. Note that, since the clean groundtruth images of SIDD and DND benchmarks are inaccessible, we upload the denoised results of all the considered methods on their online servers for testing.
@@
Qualitatively on the in-distribution SIDD validation and benchmark datasets, as shown in Fig. 9 and Fig. 10, our Condformer can not only successfully remove the noises but also keep the details well, including textures, edges and natural structures. After conditional optimization with noise prior, our Condformer could handle each noisy image in a sensor-specific way, where the denoising intensity is proportional to the corruption degree of the noisy observation. Especially under low-light imaging environments, noises prone to overwhelm the image contexts as introducing higher ISO level and indirectly enlarging the shot noises.
Furthermore, Fig. 11 and Fig. 12 show that our Condformer has excellent generalization ability as handling the out-of-distribution DND noisy images well, on both detail preservation and undesired noise removal. Especially under severe corruptions, our Condformer introduces lower distortions of color than other methods as shown in Fig. 12.
Note that, since the images have low visibility under low-light imaging environment, we employ image normalization on each denoised result for better visualization.
Table 4: Quantitative comparisons of different Transformer-based denoising methods on computational complexity. The FLOPs, Memory and Time are calculated when processing a single $\times 512$ 512 sRGB image on a NVIDIA A800 GPU.
@@
Uformer 56 50.8 343.1 2.89 0.21
Restormer 67 26.1 564.0 3.81 0.36
CAT 76 25.8 543.5 10.83 15.76
ShuffleFormer 60 50.5 344.4 2.97 0.23
GRL 30 19.8 5012.3 12.38 10.45
ART 70 25.7 542.8 3.66 0.41
@@
Table 5: Ablation study of Condformer with different noise priors on SIDD Validation dataset.
[…]
Value Location (M) (G) (s) (dB)
- - - 26.11 563.96 0.361 39.97
(0,0)(0,0)( 0 , 0 ) ($\mathbf{Q},\mathbf{K}$) 26.13 564.04 0.363 39.96
$(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ ($\mathbf{Q},\mathbf{K}$) 26.74 564.22 0.368 40.09
$(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ ($\mathbf{Q},\mathbf{K}$) 27.02 565.35 0.368 40.21
$(\sigma_{\text{s}},\sigma_{\text{r}})$ ($\mathbf{Q},\mathbf{K}$) 26.41 565.17 1.012 40.23
$(\sigma_{\text{s}},\sigma_{\text{r}})$ ($\mathbf{Q},\mathbf{K},\mathbf{V}$) 27.16 565.95 1.014 40.19
@@ 4.3.3 On computational complexity
In this section, we mainly conduct a comparison of our Condformer and other Transformer-based denoising methods on several computational complexity criteria. In detail, the number of parameters represents the model size for transmission and storage necessaries, FLOPs and Time indicate the time complexity of model, and Memory is the memory usage when running model on a GPU device, indicating the threshold level of training and inference resources.
The last three criteria are calculated when processing a single 512×512×3512$\times$ 512$\times$ 3512 × 512 × 3 sRGB image input on a NVIDIA A800 GPU. To avoid randomness, the running time is averaged on handling 100 images with torch.cuda.Event timer. As reported in Table 4, our Condformer shows relatively trade-off on the complexities, which is a resource-friendly method.
@@ 4.4 Model Analysis
Table 6: Investigation on LoNPE algorithm with different patch-sampling hyperparameters. Particularly, $\mathcal{O}$, $m/n$ and $\lambda_{\mathcal{S}}$ denote the sampling size, ratio and index, respectively. Real and synthetic scenes are considered on the public SIDD-Medium and Urban100 datasets, respectively.
LoNPE Algorithm SIDD-Medium Urban100
@@
$\times 8$ 8 555% 0.213 0.39 0.533 0.125 0.07
$\times 8$ 8 555% 0.048 0.41 0.248 0.021 0.09
$\times 8$ 8 101010% 0.040 0.52 0.233 0.020 0.11
$\times 16$ 16 101010% 0.031 0.81 0.241 0.020 0.17
$\times 16$ 16 202020% 0.036 1.40 0.325 0.031 0.28
$\times 32$ 32 101010% 0.041 1.12 0.414 0.045 0.24
LoNPE Network 0.032 0.01 0.280 0.025 0.01
@@ 4.4.1 Investigation on Condformer
Aiming at guiding the model to learn from image prior and noise prior separately, the core of our Condformer is embedding the noise prior effectively in the latent space for conditional optimization. Therefore, we conduct the ablation study in Table 5 using the same training settings as those used for the full model.
Since the core of our Condformer is constructing CondSA blocks in the latent space, we first train a baseline model which is same sa Restormer and achieves PSNR of 39.97dB on SIDD validation dataset. Subsequently, to ensure that the introduction of an auxiliary vector does not cause any performance deviation, we then build a void CondSA module by embedding a zero vector, and observe no improvement over the baseline model. However, after embedding the estimated noise prior $(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})$ from LoNPE network through a naive concatenation with the CondSA input tensor, the model achieves 0.13 PSNR gain, demonstrating the positive effect of introducing the noise prior. Furthermore, to learn an effective representation of the correlation between-in the intermediate image features and noise prior, a LFM module is designed for correlation representation of noise prior and image features. The result shows that an effective fusion module can significantly boost the contribution of the noise prior. As reported in Table 5, the 4th model achieves higher performances than the afore three models, especially gains more than 0.24dB of PSNR against the baseline.
Since noise prior is introduced as an embedding vector in CondSA, we prefer to embed it only into the query and key tensors and use the calculated attention map to enhance the value tensor. Compared to embedding the noise prior into all tensors, this design achieves higher denoising performance with lower computational complexity.
@@ 4.4.2 Investigation on LoNPE
We investigate the effects of sampling size $\mathcal{O}$ and the sampling ratio $m/n$ in Table 6, focusing on the stability of noise prior estimation by analyzing the Coefficient of Variation (CV) in scenarios where groundtruth for the noise prior is unavailable. Our analysis is conducted on the SIDD-Medium raw-domain training dataset. Specifically, since each instance in SIDD-Medium contains only two frames, we augment them using flips and rotations to compute the CV criterion. Additionally, we synthesize noisy samples by randomly generating noise from a Poisson-Gaussian distribution with $(\sigma_{\text{s}},\sigma_{\text{r}})$ set to (0.05, 0.02), (0.1, 0.05) and (0.2, 0.1). Root Mean Square Error (RMSE) criterion is calculated to evaluate the accuracy of noise prior estimation. Beside, the running time is reported in Table 6 to assess the efficiency of various sampling settings.
Our observations reveal that increasing the patch size $\mathcal{O}$ and sampling ratio $m/n$ would negatively impact the noise prior estimation as introducing more image priors, resulting in reduced stability (higher CV) and accuracy (higher RMSE). larger patches yield more precise statistical values but are more likely to include image priors, such as edges and textures, which interfere with the estimation of the noise prior. Moreover, we observe that the CV values on the SIDD-Medium dataset are significantly lower than those on Urban100, due to the lower diversity of images in SIDD-Medium. This further highlights the detrimental effect of image priors on noise prior estimation. Empirically, we select $\mathcal{O}=16\times 16$ and $m/n=10$%, as this configuration achieves a relatively better balance for estimating the noise prior across both real and synthetic scenes.
@@
Besides, to further eliminate the image prior in each patch, we introduce a local smoothness criterion $\lambda_{\mathcal{S}}$ to filter the smooth patches from the original patch pools $\{I_{i}\}^{n}_{i=1}$, which can effectively boost the accuracy of noise prior estimation as reported in Table 6. To further demonstrate the effectiveness of local smoothness criterion, we record the sampled local patches in Fig. 13. It is obvious that random sampling local patches from the raw image inevitably captures the numerous high-frequency details, which would interfere the estimation of noise prior because of the nonnegligible deviation on calculating the local luminance in Eq.(9). Instead, by employing the local smoothness criterion $\lambda_{\mathcal{S}}$, it is easy to filter the smooth patches from the original patch pools, which plays significant role on eliminating the image prior and help estimating noise prior.